In 10 years, a child will be 3 years older than twice her current age. What is the current age of the child?

Answer by stanbon(75887)

In 10 years, a child will be 3 years older than twice her current age. What is the current age of the child?
--------------------
quation:
c+10 = 3(2c)
6c = c+10
5c = 10
Child's age now is 5 yrs.
=====================
Cheers,
Stan H.
=====================